OpenStudy (anonymous):

Solve the following problem and convert to lowest terms (mixed number). 6 ⅚ – 17/18

OpenStudy (jdoe0001):

\(\bf 6\cfrac{5}{6}-\cfrac{17}{18}\implies \cfrac{6\times 6+5}{6}-\cfrac{17}{18}\implies \cfrac{41}{6}-\cfrac{17}{18}\)

OpenStudy (anonymous):

okay, fractions confuse me. So I would now subtract across or would I have to do something else?

OpenStudy (jdoe0001):

well, you'd need to get the LCD first

OpenStudy (anonymous):

okay I believe I have found the LCD What is the next step? Also, thank you for the help.

OpenStudy (anonymous):

To subtract fractions you make the denominators the same, so you times \[\frac{ 41 }{ 6 } * \frac{ 3 }{ 3 } = \frac{ 123 }{ 18}\] Then you simply subtract to top numbers to get \[\frac{ 123 }{ 18 }-\frac{ 17 }{ 18 }=\frac{ 106 }{ 18 }\] And you can simplfy that by diving both the bottom and top by 2 to get \[\frac{ 53 }{9 }\]

OpenStudy (jdoe0001):

^ that's another way to do it, you EQUATE the denominators

OpenStudy (anonymous):

Thank you for the help :)

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

@jdoe0001 always prefere multiplying to dividing when doing it mentally xD